Solve the differential equation

(z+z³) cosx dx - (z+z³) dy + (1-z²) (y-sinx) dz = 0
1
Expert's answer
2019-08-26T09:02:48-0400

Let "u=\sin x" . We have "\cos x dx= du" and "(z+z^3)du-(z+z^3)dy+(1-z^2)(y-u)dz=0" . That is "(z+z^3)(du-dy)+(1-z^2)(y-u)dz=0" .

Next, let "u-y=v" . Since "dv=du-dy" , we obtain "(z+z^3)dv+(1-z^2)(-v)dz=0."

Separate variables: "\frac{dv}{v}=\frac{1-z^2}{z+z^3}dz" We know that "\frac{dv}{v}=d\ln |v|" . Consider "\frac{1-z^2}{z+z^3}" ."z+z^3=z(z^2+1)" , so a partial fraction decomposition of "\frac{1-z^2}{z+z^3}" is "\frac{A}{z}+\frac{Bz+C}{z^2+1}=" "=\frac{A(z^2+1)+z(Bz+C)}{z(z^2+1)}" and therefore "1-z^2=A(z^2+1)+z(Bz+C)=Az^2+A+Bz^2+Cz" .

Equality of the polynomials "1-z^2" and "Az^2+A+Bz^2+Cz=(A+B)z^2+Cz+A" gives us a system of equations

"\begin{cases}
A+B=-1\\
C=0\\
A=1
\end{cases}" So, "A=1, B=-2, C=0" , and "\frac{1-z^2}{z+z^3}=\frac{1}{z}-\frac{2z}{z^2+1}" . We have "\frac{1-z^2}{z+z^3}dz=\frac{dz}{z}-\frac{2zdz}{z^2+1}="

"=d\ln |z|-\frac{dz^2}{z^2+1}=d\ln |z|-\frac{d(z^2+1)}{z^2+1}="

"=d\ln |z|-d\ln |z^2+1|=d\ln\bigl|\frac{z}{z^2+1}\bigr|" .

Since "d\ln |v| = d\ln\bigl|\frac{z}{z^2+1}\bigr|" , we have "\ln |v| = \ln\bigl|\frac{z}{z^2+1}\bigr| + D" . Take the exponent of the equation: "|v| = e^D\bigl|\frac{z}{z^2+1}\bigr|" . That is, "v = e^D\frac{z}{z^2+1}" or "v = -e^D\frac{z}{z^2+1}" . The case "v=0" also is a solution of the equation "(z+z^3)dv+(1-z^2)(-v)dz=0."

Union of the cases "v = e^D\frac{z}{z^2+1}", "v = -e^D\frac{z}{z^2+1}" and "v=0" is the case "v=E\frac{z}{z^2+1}" .

Remembering that "v=u-y" and "u=\sin x" , we have "\sin x - y = E\frac{z}{z^2+1}"


Answer: "\sin x - y = E\frac{z}{z^2+1}"